# Ahmed Heakl1  Sara Ghaboura1*  Omkar Thawakar1 Fahad Shahbaz Khan1,2   Hisham Cholakkal1   Rao Muhammad Anwer1,3   Salman Khan1,4

### v1

@@ 1.1 Quantitative Results
We present the quantitative performance of the model across benchmarks and domains. Figure 3 showcases a comprehensive performance analysis of AIN-7B across CAMEL-Bench [1] domains, comparing it with a variety of models including, GPT-4o [3], GPT-4o-mini [4], Gemini-1.5-Pro [5]), Qwen2-VL-7B [2], LlaVA-NeXt-7B [6], and Pangea-7B [7]). The plot demonstrates superior performance of AIN-7B in most domains. Table 1 shows that our AIN-7B achieves the best overall performance across all domains, compared to existing models.
@@
We conduct a comprehensive evaluation of the bilingual capabilities of AIN in Arabic and English using well-established benchmarks. To further assess AIN’s proficiency in Arabic beyond CAMEL-Bench, we evaluate its performance across all topics of ArabicMMLU [13]. As shown in Table 2, AIN outperforms in 14 out of 19 evaluated categories, achieving a significant 3.43% overall improvement compared to Qwen2-VL-7B [2]. Similarly, as presented in Table 3, AIN demonstrates strong English language capabilities on ten benchmarks that cover general VQA, mathematics, science, and visual chart interpretation.
@@ 1.3 Human Feedback
Figure 8 shows that participants predominantly favor Model 1 (AIN-7B (ours)), which received 76% of the votes. GPT-4v followed with 15%, and LLaVA garnered 9%, underscoring AIN’s significant preference among respondents (Figure 8).
Figure 8: User Model Preferences. Participant preferences for the three models in the survey, with Model 1 (AIN (ours)) receiving 76% of the votes, GPT-4v 15%, and LLaVA 9%, demonstrating AIN’s strong performance.
Figure 9: User Preferences for MSA and Local Dialects: The majority (74.3%) preferred MSA for reading and writing. An additional 11% are comfortable with MSA but favored their local dialects, while 4.3% found MSA challenging and prefer using their dialect. A further 10.5% reported difficulties unrelated to linguistic aspects.
@@
The survey results reveal that AIN outperforms human participants in several questions, demonstrating notable advantages in accuracy and adherence to response formats. For instance, in Q1, although both Model 3 and human participants provided correct answers, they failed to comply with the required response format, a criterion that AIN successfully fulfilled. Similarly, in Q9, 24.9% of participants, and in Q6, 20.5%, were unable to respond appropriately, whereas AIN provides accurate responses. In Q2, AIN exhibits superior precision by correctly identifying the shape as a disc rather than a circle, outperforming 18.6% of human respondents. Furthermore, in Q3 and Q5, participants struggle to recognize small details in the images, highlighting AIN’s ability to detect subtle features. Notably, in Q10, AIN demonstrates its capacity to solve complex reasoning tasks by extracting value beyond the visible content of the image. This highlights its ability to handle abstract problem-solving, further showcasing its comparative advantage over other models. Detailed results and comparative analyses are illustrated in Figures 10, 11, and 12.
@@ 2 Data Inspection and Selection
Our data collection comprises publicly available MSA Arabic and English datasets, with a portion specifically curated for model training. Notably, 30% of the Arabic data is authentic. To achieve scalability and address specific domain requirements, translation was utilized to complement the existing data.
@@ 2.2 Data Quality Verification and Filtering
To identify the optimal model for translation verification, we design a set of sentences reflecting common linguistic challenges in Arabic, including punctuation alignment with English, direct and semantic translation accuracy, masculine/feminine tone differentiation, and handling of diacritics. This evaluation involves 21 sentence pairs, categorized as simple sentences (Table 6), complex sentences with tone ambiguity (Table 7), and affirmative clauses with questions (Table 8).
@@ 2.5 Please sit down \<تَفَضَّل اِجلِس> No punctuation, with diacritics, masculine tone
To ensure a comprehensive quality verification process, particularly given the use of Gen-AI for translation, additional checks are necessary. These include assessing the quality of machine-generated text, its correlation with the original content, and the degree of overlap between the generated and original text. To achieve this, we employ three specialized evaluation metrics: BLEU (2-gram and 4-gram) [31] for text quality, METEOR [32] for translation correlation, and ROUGE (unigram and ROUGE-L) [33] for overlap measurement.
@@
The evaluation metrics demonstrated strong performance across multiple dimensions: BLEU scores [31] of 71.11% (2-gram) and 60.20% (4-gram) indicated high local coherence and good fluency in the translated text. The METEOR score [32] of 86.10% suggested high-quality translation with effective handling of both exact matches and linguistic variations. ROUGE metrics [33] were particularly strong, with unigram scores showing 87.80% precision and 87.30% recall, indicating excellent word-level accuracy and comprehensive content capture. Similarly, ROUGE-L scores [33] (precision: 86.20%, recall: 85.90%, F1: 85.80%) confirmed strong structural similarity between the translated and reference texts, demonstrating that the essential meaning and structure are well-preserved throughout the translation process.
@@
Following the completion of the data verification steps and toxicity filtering, our dataset, comprising 3.6M safe and curated entries, is prepared for model training.
@@ 3 Experiments
The AIN model is trained on 8 GPU nodes, each equipped with 8 NVIDIA A100 GPU cards, each with 80 GB memory. The GPUs within each node are interconnected using 8 NVLink links, ensuring high bandwidth and low latency. To facilitate efficient cross-node communication, each node is equipped with dual-port 200 Gb/sec (4×HDR) InfiniBand connections, achieving an aggregate interconnect bandwidth of 800 Gbps. This robust infrastructure was crucial in handling the computational and memory-intensive large-scale LMM training.
@@ 4 Conclusion
This work introduces AIN, an Arabic inclusive LMM, as a step toward bridging the gap in AI solutions for Arabic, a low-resource yet globally significant language. AIN is trained on a large-scale bilingual dataset with 35% of authentic Arabic data. Through rigorous evaluation, we show that AIN achieves state-of-the-art performance across a wide range of tasks, including VQA, OCR and document understanding, cultural understanding, and domain-specific applications such as medical imaging and remote sensing, surpassing even bigger and sophistcated models. AIN further demonstrates superior accuracy, contextual understanding, and human-like reasoning in MSA, as validated by extensive evaluations and human judgments. By integrating advanced data curation, robust translation pipelines, and stringent quality control, AIN provides a new state-of-the-art multimodal AI model tailored to Arabic speakers.

### v2

@@ 1.1 Quantitative Results
We present the quantitative performance of the model across benchmarks and domains. Figure 3 showcases a comprehensive performance analysis of AIN-7B across CAMEL-Bench [1] domains, comparing it with a variety of models including, GPT-4o [3], GPT-4o-mini [4], Gemini-1.5-Pro [5], Qwen2-VL-7B [2], LlaVA-NeXt-7B [6], and Pangea-7B [7]. The plot demonstrates the superior performance of AIN-7B in most domains. Table 1 shows that our AIN-7B achieves the best overall performance across all domains, compared to existing models.
@@
We conduct a comprehensive evaluation of the bilingual capabilities of AIN in Arabic and English using well-established benchmarks. To further assess AIN’s proficiency in Arabic beyond CAMEL-Bench, we evaluate its performance across all topics of ArabicMMLU [13]. As shown in Table 2, AIN outperforms in 14 out of 19 evaluated categories, achieving a significant 3% overall improvement compared to Qwen2-VL-7B [2]. Similarly, as presented in Table 3, AIN demonstrates strong English language capabilities on ten benchmarks that cover general VQA, mathematics, science, and visual chart interpretation.
@@ 1.3 Human Feedback
Figure 8 shows that participants predominantly favor Model 1 (AIN-7B (ours)), which received 76% of the votes. GPT-4o followed with 15%, and LLaVA garnered 9%, underscoring AIN’s significant preference among respondents.
Figure 8: User Model Preferences. Participant preferences for the three models in the survey, with Model 1 (AIN (ours)) receiving 76% of the votes, GPT-4o 15%, and LLaVA 9%, demonstrating AIN’s strong performance.
Figure 9: User Preferences for MSA and Local Dialects: The majority (74.3%) preferred MSA for reading and writing. An additional 11% are comfortable with MSA but favored their local dialects, while 4.3% found MSA challenging and preferred using their dialect. A further 10.5% reported difficulties unrelated to linguistic aspects.
@@
The survey results reveal that AIN outperforms human participants in several questions, demonstrating notable advantages in accuracy and adherence to the response formats. For instance, in Q1, although both Model 3 and human participants provided correct answers, they failed to comply with the required response format, a criterion that AIN successfully fulfilled. Similarly, in Q9, 24.9% of participants, and in Q6, 20.5%, were unable to respond appropriately, whereas AIN provides accurate responses. In Q2, AIN exhibits superior precision by correctly identifying the shape as a disc rather than a circle, outperforming 18.6% of human respondents. Furthermore, in Q3 and Q5, participants struggle to recognize small details in the images, highlighting AIN’s ability to detect subtle features. Notably, in Q10, AIN demonstrates its capacity to solve complex reasoning tasks by extracting value beyond the visible content of the image. This highlights its ability to handle abstract problem-solving, further showcasing its comparative advantage over other models. Detailed results and comparative analyses are illustrated in Figures 10, 11, and 12.
@@ 2 Data Inspection and Selection
Our data collection comprises publicly available MSA Arabic and English datasets, with a portion specifically curated for model training. Notably, 35% of the Arabic data is authentic. To achieve scalability and address specific domain requirements, translation was utilized to complement the existing data.
@@ 2.2 Data Quality Verification and Filtering
To identify the optimal model for translation verification, we design a set of sentences reflecting common linguistic challenges in Arabic, including punctuation alignment with English, direct and semantic translation accuracy, masculine/feminine tone differentiation, and handling of diacritics. This evaluation involves 21 sentence pairs, categorized as simple sentences (Table 6), complex sentences with tone ambiguity (Table 7), and affirmative clauses with question (Table 8).
@@ 2.5 Please sit down \<تَفَضَّل اِجلِس> No punctuation, with diacritics, masculine tone
To ensure a comprehensive quality verification process, particularly given the use of generative AI for translation, additional checks are necessary. These include assessing the quality of machine-generated text, its correlation with the original content, and the degree of overlap between the generated and original text. To achieve this, we employ three specialized evaluation metrics: BLEU (2-gram and 4-gram) [31] for text quality, METEOR [32] for translation correlation, and ROUGE (unigram and ROUGE-L) [33] for overlap measurement.
@@
The evaluation metrics demonstrated strong performance across multiple dimensions: BLEU scores [31] of 71.11% (2-gram) and 60.20% (4-gram) indicated high local coherence and good fluency in the translated text. The METEOR score [32] of 86.10% suggested high-quality translation with effective handling of both exact matches and linguistic variations. ROUGE metrics [33] were particularly strong, with unigram scores showing 87.80% precision and 87.30% recall, indicating excellent word-level accuracy and comprehensive content capture. Similarly, ROUGE-L scores [33] (precision: 86.20%, recall: 85.90%, F1: 85.80%) confirmed strong structural similarity between the translated and reference texts, demonstrating that the essential meaning and structure are well-preserved throughout the translation process (Table 9).
@@
Following the completion of the data verification steps and toxicity filtering, our dataset, comprising 3.6 million safe and curated entries, is prepared for model training.
@@ 3 Experiments
The AIN model is trained on 8 GPU nodes, each equipped with 8 NVIDIA A100 GPU cards, each with 80 GB memory. The GPUs within each node are interconnected using 8 NVLink links, ensuring high bandwidth and low latency. To facilitate efficient cross-node communication, each node is equipped with dual-port 200 Gbps (4×HDR) InfiniBand connections, achieving an aggregate interconnect bandwidth of 800 Gbps. This robust infrastructure was crucial in handling the computational and memory-intensive large-scale LMM training.
@@ 4 Conclusion
This work introduces AIN, an Arabic-inclusive LMM, as a step toward bridging the gap in AI solutions for Arabic, a low-resource yet globally significant language. AIN is trained on a large-scale bilingual dataset with 35% of authentic Arabic data. Through rigorous evaluation, we show that AIN achieves state-of-the-art performance across a wide range of tasks, including VQA, OCR and document understanding, cultural understanding, and domain-specific applications such as medical imaging and remote sensing, surpassing even bigger and more sophisticated models. AIN further demonstrates superior accuracy, contextual understanding, and human-like reasoning in MSA, as validated by extensive evaluations and human judgments. By integrating advanced data curation, robust translation pipelines, and stringent quality control, AIN provides a new state-of-the-art multimodal AI model tailored to Arabic speakers.